Fermi National Laboratory


Accelerator Update for February 7, 2005 - February 9, 2005

Monday February 7
The day shift began with Operations monitoring store 3969 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

Operations established beam to the Meson MCenter experiment at 8:19 AM.

The Neutron Therapy Facility (NTF) requested beam at 9:26 AM.

At 11:35 AM, the Tevatron (TeV) suffered a sector A1 and E1 quench. Experts investigated a possible RF and damper problem.

Operators made the Collider Detector at Fermilab (CDF) ready for access at 12:45 PM. They came out of access at 1:24 PM.

Operations put the TeV through a dry squeeze at 2:49 PM and then began a study period.

An expert reported at 3:39 PM that the TeV dampers were okay.

At 5:29 PM, the Main Injector (MI) ramp tripped off on a ground fault indication. Operators reset the system and it tripped again. Operators began a MI hipot.

Operations reported at 6:53 PM that they had completed all hipots with no indication of any ground faults, but, unfortunately, the MI ramp continued to trip.

At 7:42 PM, a MI expert narrowed the cause of the MI trips to a power supply at MI-10. A power supply expert will investigate further.

Tuesday February 8
The midnight shift began with Operations monitoring the stash, with MI and power supply experts searching for a MI ground fault, and with MiniBooNE taking beam. With no beam in MI, neither Pbar nor Switchyard (SY) received beam.

MI and power supply experts reported at 2:04 AM that they found a polarity switch problem in a MI10L power supply. They completed their repairs at 2:47 AM.

Operations reported at 3:26 AM that MI was ramping and that they had verified stacking.

Operations put the TeV into shot setup at 3:46 AM.

At 4:39 AM, Operations reported that a power supply (VTF17) wasn't ramping. Operators discovered the supply's control tables were blank and restored their values, but the fault wouldn't clear. At 5:25 AM, operators reported that they had changed the control cards, but that didn't work either.

Operations reported at 5:47 AM that they left VTF17 off and compensated for it by raising the current on another correction magnet power supply (F17B3).

At 5:55 AM, a TeV sector F2 wet engine came into alarm on overtemp. Operators informed the cryo system technicians at CHL, who investigated and found a bad fan.

Operations established store 3972 at 9:16 AM with an initial luminosity of 78.98E30.

Operations established beam to the Meson MCenter experiment at 10:06 AM.

A power supply expert completed repairs on VTF17 at 10:10 AM.

Operations reported at 10:34 AM that after the VTF17 repairs were completed operators were able to send good beam to MTest and MCenter.

Operators turned off MiniBooNE in preparation of this afternoon's access into MI and Booster.

Operations turned off beam at 12:15 PM and made the necessary preparations for accesses into MI, Booster, and Linac.

A Linac expert switched to the H- Source at 2:22 PM. The I- Source needed its generator rebuilt.

MI experts reported by 3:46 PM that the RF and kicker work was finished.

The Booster RF was up and running by 4:28 PM.

At 5:08 PM, Operations reported that the VCB for MI-50 would not close. It closed for a second and then a breaker tripped.

At 7:24 PM, Operations reported that the MI-50 quadrupole power supply VCB switch gear cabinet suffered a failure. MI is off. Experts have been notified. There will be no beam to Pbar or to the Meson experiments.

Wednesday February 9
The midnight shift began with Operations monitoring store 3972, with MI off, with the Pbar stack at 40.3E10 and being monitored, and with MiniBooNE taking beam.

At 5:05 AM, a Linac RF station (KRF2) tripped off and wouldn't reset. This shut off MiniBooNE beam.

The Plan for Wednesday
The plan for today is to keep the TeV store in as long as possible. MI is down until the problem with the switch gear is understood. This may take two to three days.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/9/2005   email Fermilab